OpenStudy (anonymous):

is the are enclosed between y=3x and y=x^2 a 9/2 b 51/4

OpenStudy (anonymous):

area

OpenStudy (anonymous):

iv got \[\frac{ 3 }{ 2 }x ^{2}-\frac{ 1 }{ 3 }x ^{3} 0 and 3\]

OpenStudy (anonymous):



OpenStudy (anonymous):

what is the integral that you should calculate ?

OpenStudy (anonymous):

well that's good

OpenStudy (anonymous):

what you wrote up there

OpenStudy (anonymous):

\[\int\limits_{0}^{3} 3x-x ^{2}dx\]

OpenStudy (anonymous):

what do i do after i have reached that point?

OpenStudy (anonymous):

yes that's good.. so what is your answer

OpenStudy (anonymous):